OpenStudy (ashy98):

2 + a + b = a + 2 + b is an example of which algebraic property? A.) Distributive Property B.) Associative Property of Addition C.) Commutative Property of Addition D.) Symmetric Property

OpenStudy (dangerousjesse):

Do you know what all of these properties are?

OpenStudy (ashy98):

the truth no!

OpenStudy (dangerousjesse):

Okay :) An example of the distributive property is \(\large 2(x+y)=2x+2y\)

OpenStudy (dangerousjesse):

So the equation isn't distributive.

OpenStudy (dangerousjesse):

The associative property of addition is just where you move around parentheses in an equation, like \(\large (x+y)+z=x+(y+z)\).

OpenStudy (dangerousjesse):

There are no parentheses in the equation, so it's not associative.

OpenStudy (ashy98):

Symmetric Property?

OpenStudy (dangerousjesse):

Commutative is just showing that you can mix up the numbers, like \(\large x+y+z=z+x+y\).

OpenStudy (dangerousjesse):

And symmetric shows a reflection as the sum. \(\large x+y+z=z+y+x\)

OpenStudy (dangerousjesse):

So what will be your answer?

OpenStudy (ashy98):

A

OpenStudy (dangerousjesse):

How is the equation distributing the variables amongst each other?

OpenStudy (ashy98):

C.) Commutative Property of Addition

OpenStudy (dangerousjesse):

Yes! :)

OpenStudy (ashy98):

yay thank you soo muchh !!!(: